# DiffMesh: A Motion-aware Diffusion-like Framework for Human Mesh Recovery from Videos

### v3

@@ Abstract
Human mesh recovery (HMR) provides rich human body information for various real-world applications. While image-based HMR methods have achieved impressive results, they often struggle to recover humans in dynamic scenarios, leading to temporal inconsistencies and non-smooth 3D motion predictions due to the absence of human motion. In contrast, video-based approaches leverage temporal information to mitigate this issue. In this paper, we present DiffMesh, an innovative motion-aware Diffusion-like framework for video-based HMR. DiffMesh establishes a bridge between diffusion models and human motion, efficiently generating accurate and smooth output mesh sequences by incorporating human motion within the forward process and reverse process in the diffusion model. Extensive experiments are conducted on the widely used datasets (Human3.6M [15] and 3DPW [48]), which demonstrate the effectiveness and efficiency of our DiffMesh. Visual comparisons in real-world scenarios further highlight DiffMesh’s suitability for practical applications. The project webpage is https://zczcwh.github.io/diffmesh_page/.
@@ 1 Introduction
Recently, diffusion models have garnered substantial interest within the computer vision community. By learning to reverse the diffusion process in which noise has been added in finite successive steps, diffusion models can generate samples that match a specified data distribution corresponding to the provided dataset. As generative-based methods, diffusion models excel in tasks such as motion synthesis and generation [45, 18, 28]. The quality and diversity of generated outputs have witnessed significant improvements by bridging the gap between uncertain and determinate distributions during the denoising process. This inspires us to consider the diffusion model as a promising solution for recovering high-quality human mesh from input data fraught with depth ambiguity.
While it is straightforward to apply the diffusion model for image-based HMR such as in [11], effective utilization of the diffusion model for video-based HMR remains a formidable challenge. As previous video-based HMR methods have not yet incorporated diffusion models into their frameworks, we provide a summary of diffusion-based methods related to human pose and mesh recovery (please refer to Sec. 2.2 for further elaboration). We then organize these methods into two intuitive baseline approaches that can be applied for video-based HMR, as illustrated in Fig. 1 (a) and (b). Assuming a video input with $f$ frames in Baseline 1, the diffusion model is applied individually to recover the mesh of the single frame (usually would be the center frame). More details can be found in Sec. 3.2. However, this approach suffers from significant computational burdens as the total number of denoising steps amounts to $f\times N$ to recover the mesh sequence of $f$ frames, with $N$ representing the number of denoising steps in the diffusion model. Moreover, the diffusion model does not account for motion smoothness in dynamic sequences, which may result in potentially temporal inconsistent mesh predictions. To mitigate computational costs and enhance inference speed, we introduce Baseline 2. The features for ground truth mesh vertices of each frame $x_{i}$, where $i\in\left\{1,\dots f\right\}$ are concatenated to unified features $x_{unified}$ before the forward diffusion process. Subsequently, the output features are partitioned back into individual frames $x_{i}$ after the denoising process. While this strategy effectively reduces the number of steps from $f\times N$ to $N$, it is important to note that motion smoothness in dynamics is still not considered during the forward and reverse process since features across frames are simply concatenated. This limitation can potentially lead to non-smooth motion predictions.
@@
To tackle the aforementioned efficiency and motion smoothness challenges, we introduce DiffMesh, a novel motion-aware diffusion-like framework as illustrated in Fig. 1 (c). DiffMesh is designed specifically for video-based HMR. Unlike conventional diffusion models, which necessitate the addition of Gaussian noise in consecutive steps during the forward diffusion process, as depicted in Fig. 2 (a), DiffMesh adopts a distinct approach. We consider the forward motion influence over a brief time interval as shown in Fig. 2 (b), encompassing the features (embedded in the high-dimensional space) of adjacent frames $x_{i}$ to $x_{i+1}$, similar to the mechanism of introducing noise. Importantly, our approach does not conclude the forward motion at the last frame $x_{f}$, where $f$ is the total number of frames. Instead, we assume that motion culminates in an initial distribution represented by the static state, which is similar to pure Gaussian noise in the conventional diffusion model, after additional forward steps. During the reverse process, the diffusion model decodes the previous feature $x_{i-1}$ based on $x_{i}$ and the corresponding conditional feature $c_{i}$ from the input frame. Consequently, the total number of steps remains consistent at $N$, as illustrated in Fig. 1 (b).
Contribution: The key innovation of DiffMesh lies in its capacity to consider the inherent motion patterns within the diffusion model efficiently and effectively. This consideration notably enhances motion smoothness and temporal consistency throughout the output mesh sequence, achieved through successive steps across frames. Incorporating our diffusion framework, we design a new two-stream transformer architecture to decode motion features. Our DiffMesh not only achieves new state-of-the-art results on the Human3.6M [15] and 3DPW [48] datasets, but also demonstrates its superiority in handling in-the-wild scenarios compared to previous methods. These advantages make DiffMesh more viable for deployment in real-world applications for generating accurate and smooth human mesh sequences efficiently.
@@ 2 Related Work
Since human mesh recovery is a popular topic in computer vision, here we focus on the most relevant works to our methods. We refer readers to the recent and comprehensive HMR survey [46] for more details.
@@ 2.1 Video-based HMR
Similar to the image-based methods  [30, 57, 62, 53, 22, 27, 61], video-based methods [17, 32, 44, 55, 25, 55, 42, 54, 52] for HMR also utilize parametric human models like SMPL [31]. VIBE [20] leverages the AMASS dataset [33] to discriminate between real human motions and those produced by its temporal encoder. MEVA [32] estimates the coarse 3D human motion using a variational motion estimator, then refines the motion by a motion residual regressor. TCMR [6] utilizes GRU-based temporal encoders to forecast the current frame from the past and future frames, which enhances the temporal consistency. Similar to [6], MPS-Net [50] captures the motion continuity dependencies by the attention mechanism. MotionBERT [64] proposes a pertaining stage to recover 3D motion from noise 2D observations, achieving state-of-the-art performance.
@@ 2.2 Diffusion Generative Models
DiffPose [10] Video sequence Baseline 1
@@
Table 1: Comparison with previous human pose and mesh-related methods utilizing diffusion models.
Distinction of Our Method: Previous methods for video-based HMR have not integrated diffusion models into their frameworks. While there are diffusion-based methods that can output 3D human pose or human mesh, they are not directly suitable for the video-based HMR task and require adaptations. We organize them into two intuitive baseline approaches. As depicted in Table 1, approaches similar to Baseline 1, such as [10, 11, 12], require $f\times N$ denoising steps to produce output for $f$ frames. Conversely, methods similar to Baseline 2, such as [63, 41, 28], reduce the total denoising steps from $f\times N$ to $N$. However, these methods do not effectively incorporate the consideration of human motion within the sequence during both the forward process and the reverse process, which may potentially result in temporal inconsistencies and non-smooth 3D motion predictions. In contrast, our approach treats motion patterns across frames as the inherent noise in the forward process, eliminating the need for adding Gaussian noise during forward steps as required in Baseline 1 and Baseline 2. This unique motion-aware design in diffusion enables our approach to decode human motion patterns during reverse processes within $N$ steps, setting it apart from conventional diffusion-based methods.
@@ 3.1 Preliminary
$\small p(x_{t}|x_{0}=\mathcal{N}(x_{t};\sqrt{\hat{\beta_{t}}}\cdot x_{0},(1-\hat{\beta_{t}})\cdot\textbf{I}),\vspace{-10pt}$ (2)
@@
where $\theta$ denotes model parameters, and the mean $\mu_{\theta}(x_{t},t)$ and variance $\Sigma_{\theta}(x_{t},t)$ are parameterized by deep neural networks $\epsilon_{\theta}$.
@@ 3.2 Diffusion-based Baselines
Figure 4: The architecture of DiffMesh: Our framework takes input sequence with $f$ frames, with the objective of outputting a human mesh sequence consisting of $f$ frames. We model the forward human motion across frames similar to the mechanism of introducing noise in the forward process. We assume that human motion will eventually reach a static state, which is represented by the mesh template state $x_{end}$. Thus, additional $N-f+1$ steps are necessary from $x_{f}$ to reach the static state $x_{end}$. Consequently, we utilize a transformer-based diffusion model to sequentially produce the decoded features during the reverse process. The final human mesh sequence is returned by a mesh head using SMPL [31] human body model. The structure of DN is illustrated in Fig. 5.
Given that previous video-based HMR methods have not integrated diffusion models into their frameworks, we incline to explore the performance of simply incorporating diffusion models in the context of video-based HMR.
@@
Baseline 1: As illustrated in Fig. 3 (a), we initially apply the vanilla diffusion model at the frame level. For each $i_{th}$ frame, we employ a backbone to extract the conditional feature $c_{i}$ where $i\in\left\{1,\dots,f\right\}$ and $f$ representing the total number of frames. We then proceed with the standard forward diffusion process, introducing noise to the ground truth mesh vertices $x_{i}$. After $N$ steps, the feature $x_{i}^{end}$ converges to a Gaussian Noise. Subsequently, a transformer-based diffusion model (details are included in the supplementary Sec. B) is employed for the denoising process. The denoised feature $x^{\prime}_{i}$ is obtained after $N$ denoising steps. Finally, the human mesh of $i_{th}$ frame is returned by the human mesh head. By iterating this procedure $f$ times, we generate the final human mesh sequence consisting of $f$ frames. However, despite its simplicity and straightforwardness, Baseline 1 exhibits a significant drawback: the total number of steps is $f\times N$, resulting in considerable computational overhead and slow inference speed. Furthermore, it fails to address motion smoothness within the diffusion model, potentially leading to temporal inconsistencies in mesh predictions.
@@
Baseline 2: To mitigate computational costs and enhance inference speed, we introduce Baseline 2 for video-based HMR. Motivated by recent works [59, 18, 28] that leverage diffusion models for motion generation, we adopt a similar concatenation and separation strategy, as illustrated in Fig. 3 (b). In this approach, we concatenate the ground-truth mesh vertices of each frame into a unified feature $x_{unified}=cat(x_{1},x_{2},\dots,x_{f})$, where $f$ represents the total number of frames. Similarly, we create a concatenated conditional feature $c_{unified}$ by combining the image features from each frame. Subsequently, we apply the forward diffusion process by introducing noise for $N$ steps, which is analogous to Baseline 1. After $N$ steps, $x_{unified}^{end}$ also converges to a Gaussian Noise state. Following this, we perform $N$ denoising steps utilizing a transformer-based diffusion model (details are included in the supplementary Sec. B). The denoised feature $x^{\prime}_{unified}$ is then partitioned into $x^{\prime}_{1},x^{\prime}_{2},\dots,x^{\prime}_{f}$. Finally, the mesh sequence is obtained through the mesh head.
@@
Despite the substantial reduction in the number of steps, achieved by the concatenation and separation strategy from $f\times N$ to $N$, it is important to note that this strategy does not effectively consider motion patterns across frames due to simple concatenation, which may result in temporal inconsistencies and non-smooth 3D mesh predictions.
@@ 3.3 DiffMesh
To address the previously mentioned limitations in our baselines, we present DiffMesh, a novel solution for video-based HMR. The overall architecture is depicted in Fig. 4. Our framework takes input sequence with $f$ frames, with the objective of outputting a human mesh sequence consisting of $f$ frames. Additionally, we extract the conditional feature $c_{i}$ using the backbone, incorporating a conditional features generation block (more details are provided in the supplementary Sec. B). In contrast to Baseline 1 and Baseline 2, DiffMesh employs a distinct approach that accounts for the forward and backward motion within the diffusion model.
Forward Process: We make the assumption that human motion consistently influences the human mesh vertices (in the embedded high-dimensional space) across frames, denoted as $x_{i}$ to $x_{i+1}$, where $i\in\left\{1,\dots f-1\right\}$. This motion pattern can be conceptualized as a specific noise introduced to the mesh vertices feature of the preceding frame:
@@
where $m_{i}$ represents the motion pattern from $i_{th}$ frame to $(i+1)_{th}$ frame. Consequently, $x_{1}$ can be gradually diffused into $x_{f}$ after $f-1$ forward diffusion steps. Importantly, achieving $x_{f}$ does not denote the conclusion of the forward diffusion process. We assume that human motion will eventually reach a static state $x_{end}$, which is an initial distribution (similar to the Gaussian noise in the conventional diffusion model). In our work, we select the template mesh acquired by averaging ground-truth meshes [30] as the initial distribution. To arrive at this final static state $x_{end}$, DiffMesh requires $N-f+1$ additional forward diffusion steps. Therefore, the total number of steps remains $N$.
@@
Reverse Process: Our objective during the reverse process is to obtain output features $x^{\prime}_{1}\cdots x^{\prime}_{f}$. Initiating from the static state $x_{end}$, we design a transformer-based diffusion model, as depicted in Fig. 4, to sequentially produce decoded features. After $N-f+1$ steps, our diffusion model yields $x^{\prime}_{f}$, which serves as the target feature for the final frame $f$. Iteratively, we generate the output features $x^{\prime}_{i}$ where $i\in\left\{f-1,\dots 1\right\}$ over $f-1$ steps. The total number of steps remains $N$. In contrast to Baseline 2, DiffMesh demonstrates the capability to decode specific motion patterns within the same number of steps, while generating more precise and smooth output (verified in Sec. 4.4).
Network Design: In contrast to conventional diffusion-based methods [14, 38, 37] for image synthesis tasks that often rely on UNet [39] or vanilla transformer [47] backbones, our approach introduces a two-stream transformer design deeply integrated within our framework, as illustrated in Fig. 5. While conventional diffusion-based methods [14, 38, 37] only estimate noise during steps, our two-stream network predicts the motion features and the previous conditional feature during each step as depicted in Fig. 5 (c). This architecture significantly enhances information integration through a two-stream process. Initially, it captures mesh and condition dependencies separately using two self-attention blocks. Then, these captured dependencies are merged by a cross-attention block efficiently to decode motion features. An auxiliary loss is computed by measuring the difference between the predicted $\hat{c}_{i-1}$ and the ground truth $c_{i-1}$, which contributes to training a better transformer-based network during the reverse process.
Given the input feature $x^{\prime}_{i}$ and corresponding conditional feature $c_{i}$ at each step, our network produces the predicted motion $m_{i-1}$ as well as the predicted previous conditional feature $\hat{c}_{i-1}$. The feature of static state (mesh template) is denoted as $x_{end}$. Similar to Equation 5, the mesh features $x^{\prime}_{i-1}$ in DiffMesh can be computed as
$\small x^{\prime}_{i-1}=\frac{1}{\sqrt{\beta_{i}}}(x^{\prime}_{i}-\frac{1-\beta_{i}}{\sqrt{\hat{1-\beta_{i}}}}{\color[rgb]{1,.5,0}m_{i-1}})+\sigma_{t}{\color[rgb]{1,.5,0}x_{end}},$ (7)
@@
Figure 5: (a) The architecture of our two-stream transformer network. Given the input feature $x^{\prime}_{i}$ and corresponding conditional feature $c_{i}$, the diffusion model generates the predicted motion $m_{i-1}$ and the predicted previous conditional feature $\hat{c}_{i-1}$. (b) The self-attention block in our diffusion model. (c) The cross-attention block in our diffusion model. The notation $\oplus$ represents element-wise addition.

### v4

# DiffMesh: A Motion-aware Diffusion Framework for Human Mesh Recovery from Videos

@@ Abstract
Human mesh recovery (HMR) provides rich human body information for various real-world applications such as gaming, human-computer interaction, and virtual reality. While image-based HMR methods have achieved impressive results, they often struggle to recover humans in dynamic scenarios, leading to temporal inconsistencies and non-smooth 3D motion predictions due to the absence of human motion. In contrast, video-based approaches leverage temporal information to mitigate this issue. In this paper, we present DiffMesh, an innovative motion-aware diffusion framework for video-based HMR. DiffMesh establishes a bridge between diffusion models and human motion, efficiently generating accurate and smooth output mesh sequences by incorporating human motion within the forward process and reverse process in the diffusion model. Extensive experiments are conducted on the widely used datasets (Human3.6M [16] and 3DPW [51]), which demonstrate the effectiveness and efficiency of our DiffMesh. Visual comparisons in real-world scenarios further highlight DiffMesh’s suitability for practical applications. The code will be publicly available.
@@ 1 Introduction
Recently, diffusion models have garnered substantial interest within the computer vision community. By learning to reverse the diffusion process in which noise has been added in successive steps, diffusion models can generate samples that match a specified data distribution corresponding to the provided dataset. As generative-based methods, diffusion models excel in tasks such as motion synthesis and generation [48, 20, 27]. The quality and diversity of generated outputs have witnessed significant improvements by bridging the gap between uncertain and determinate distributions. This inspires us to consider the diffusion model as a promising solution for recovering high-quality human mesh from input data fraught with depth ambiguity.
While it is straightforward to apply the diffusion model for image-based HMR such as in [11], effective utilization of the diffusion model for video-based HMR remains a formidable challenge. As previous video-based HMR methods have not yet incorporated diffusion models into their frameworks, we provide a summary of diffusion-based methods related to human pose and mesh recovery (please refer to Sec. 2.2 for further elaboration). We then organize these methods into two intuitive baseline approaches that can be applied for video-based HMR, as illustrated in Fig. 1 (a) and (b). Assuming a video input with $f$ frames in Baseline 1, the diffusion model is applied individually to recover the mesh of the single frame (usually would be the center frame). More details can be found in Sec. 3.2. However, this approach suffers from significant computational burdens as the total number of denoising steps amounts to $f\times N$ to recover the mesh sequence of $f$ frames, with $N$ representing the number of denoising steps in the diffusion model. Moreover, the diffusion model does not account for motion smoothness in dynamic sequences, which may result in potentially temporal inconsistent mesh predictions. To mitigate computational costs and enhance inference speed, we introduce Baseline 2. The features for ground truth mesh vertices of each frame $x_{i}$, where $i\in\left\{1,\dots f\right\}$ are concatenated to unified features $x_{unified}$ before the forward diffusion process. Subsequently, the output features are partitioned back into individual frames $x_{i}$ after the denoising process. While this strategy effectively reduces the steps from $f\times N$ to $N$, it is important to note that motion smoothness in dynamics is still not considered during the forward and reverse process since features across frames are simply concatenated. This limitation can potentially lead to non-smooth motion predictions.
@@
To tackle the aforementioned efficiency and motion smoothness challenges when utilizing the diffusion model, we introduce DiffMesh, a novel motion-aware diffusion framework as illustrated in Fig. 1 (c). DiffMesh is designed specifically for video-based HMR. Unlike conventional diffusion models, which necessitate the addition of Gaussian noise from the input image during the forward diffusion process, as depicted in Fig. 2 (a), DiffMesh adopts a distinct approach. We consider the forward motion influence over a brief time interval as shown in Fig. 2 (b), encompassing the features (embedded in the high-dimensional space) of adjacent frames $x_{i}$ to $x_{i+1}$, similar to the mechanism of introducing noise. Importantly, our approach does not conclude the forward motion at the last frame $x_{f}$, where $f$ is the total number of frames. Instead, we assume that the initial distribution can be reached after continuously adding forward motion by finite diffusion steps, which is similar to the Gaussian distribution in the conventional diffusion model. During the reverse process, the diffusion model decodes the previous feature $x_{i-1}$ based on $x_{i}$ and the corresponding conditional feature $c_{i}$ from the input frame. Consequently, the total number of steps remains consistent at $N$, as illustrated in Fig. 1 (b).
Contribution: The key innovation of DiffMesh lies in its capacity to consider the inherent motion patterns within the diffusion model efficiently and effectively. This consideration notably enhances motion smoothness and temporal consistency throughout the output mesh sequence, achieved through successive steps across frames. Incorporating our diffusion framework, we design a two-stream transformer architecture to decode motion features. Our DiffMesh not only achieves new state-of-the-art results on the Human3.6M [16] and 3DPW [51] datasets, but also demonstrates its superiority in handling in-the-wild scenarios compared to previous methods. These advantages make DiffMesh more viable for deployment in real-world applications for generating accurate and smooth human mesh sequences efficiently.
@@ 2 Related Work
Since human mesh recovery is a popular topic in computer vision, here we focus on the most relevant works to our methods. More references can be found in the supplementary Sec. 2. We Also refer readers to the recent and comprehensive HMR survey [49] and diffusion model survey [55] for more details.
@@ 2.1 Video-based HMR
Image-based HMR methods  [32, 68, 25, 66, 62, 30, 12, 53, 19] have achieved remarkable progress in accuracy. However, when applied to video sequences, these image-based methods suffer from severe motion jitter due to frame-by-frame reconstruction, making them unsuitable for practical use. To address this issue, video-based methods [18, 34, 47, 60, 26, 60, 45, 59, 56] utilize temporal information to enhance accurate and temporally consistent human mesh from video frames. VIBE [22] leverages the AMASS dataset [35] to discriminate between real human motions and those produced by its temporal encoder. MEVA [34] estimates the coarse 3D human motion using a variational motion estimator, then refines the motion by a motion residual regressor. TCMR [5] utilizes GRU-based temporal encoders to forecast the current frame from the past and future frames, which enhances the temporal consistency. Similar to [5], MPS-Net [54] captures the motion continuity dependencies by the attention mechanism. MotionBERT [70] proposes a pertaining stage to recover 3D motion from noise 2D observations, achieving state-of-the-art performance.
@@ 2.2 Diffusion Generative Models
DiffPose [13] 2D Human Pose Sequence Baseline 1
@@
Table 1: Comparison with previous diffusion-based human pose and mesh methods.
Distinction of Our Method: Previous methods for video-based HMR have not integrated diffusion models into their frameworks. While there are diffusion-based methods that can output 3D human pose or human mesh, they are not directly suitable for the video-based HMR task and require adaptations. We organize them into two intuitive baseline approaches. As depicted in Table 1, approaches similar to Baseline 1, such as [10, 11, 13], require $f\times N$ denoising steps to produce output for $f$ frames. Conversely, methods similar to Baseline 2, such as [69, 44, 27], reduce the total denoising steps from $f\times N$ to $N$. However, these methods do not effectively incorporate the consideration of human motion within the sequence during both the forward process and the reverse process, which may potentially result in temporal inconsistencies and non-smooth 3D motion predictions. In contrast, our approach treats motion patterns across frames as the inherent noise in the forward process. This unique motion-aware design in diffusion enables our approach to decode human motion patterns during reverse processes within $N$ steps, setting it apart from conventional diffusion-based methods.
@@ 3.1 Preliminary
$\small p(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\hat{\beta_{t}}}\cdot x_{0},(1-\hat{\beta_{t}})\cdot\textbf{I}),\vspace{-10pt}$ (2)
@@
where $\theta$ denotes model parameters, and the mean $\mu_{\theta}(x_{t},t)$ and covariance $\Sigma_{\theta}(x_{t},t)$ are parameterized by deep neural networks $\epsilon_{\theta}$.
@@ 3.2 Diffusion-based Baselines
Figure 3: Transformation between human mesh and corresponding mesh feature in the latent space.
Given that previous video-based HMR methods have not integrated diffusion models into their frameworks, we incline to explore the performance of simply incorporating diffusion models in the context of video-based HMR. The forward and reverse processes are conducted in a latent space, the transformation between human mesh and the corresponding mesh feature in this latent space is shown in Fig. 3.
@@
Baseline 1: As illustrated in Fig. 4 (a), we initially apply the vanilla diffusion model at the frame level. For each $i_{th}$ frame, we employ a backbone to extract the conditional feature $c_{i}$ where $i\in\left\{1,\dots,f\right\}$ and $f$ representing the total number of frames. We then proceed with the standard forward diffusion process, introducing noise to the ground truth mesh vertices $x_{i}$. After $N$ steps, the feature $x_{i}^{end}$ converges to a Gaussian Noise. Subsequently, a transformer-based diffusion model (details are included in the supplementary Sec. 2) is employed for the denoising process. The denoised feature $x^{\prime}_{i}$ is obtained after $N$ denoising steps. Finally, the human mesh of $i_{th}$ frame is returned by the human mesh head. By iterating this procedure $f$ times, we generate the final human mesh sequence consisting of $f$ frames. However, despite its simplicity and straightforwardness, Baseline 1 exhibits a significant drawback: the total number of steps is $f\times N$, resulting in considerable computational overhead and slow inference speed. Furthermore, it fails to address motion smoothness within the diffusion model, potentially leading to temporal inconsistencies in mesh predictions.
@@
Baseline 2: To mitigate computational costs and enhance inference speed, we introduce Baseline 2 for video-based HMR. Motivated by recent works [65, 20, 27] that leverage diffusion models for motion generation, we adopt a similar concatenation and separation strategy, as illustrated in Fig. 4 (b). In this approach, we concatenate the ground-truth mesh vertices of each frame into a unified feature $x_{unified}=cat(x_{1},x_{2},\dots,x_{f})$, where $f$ represents the total number of frames. Similarly, we create a concatenated conditional feature $c_{unified}$ by combining the image features from each frame. Subsequently, we apply the forward diffusion process by introducing noise for $N$ steps, which is analogous to Baseline 1. After $N$ steps, $x_{unified}^{end}$ also converges to a Gaussian Noise state. Following this, we perform $N$ denoising steps utilizing a transformer-based diffusion model (details are included in the supplementary Sec. 2). The denoised feature $x^{\prime}_{unified}$ is then partitioned into $x^{\prime}_{1},x^{\prime}_{2},\dots,x^{\prime}_{f}$. Finally, the mesh sequence is obtained through the mesh head.
Figure 5: The architecture of DiffMesh: Our framework takes input sequence with $f$ frames, with the objective of outputting a human mesh sequence consisting of $f$ frames. We model the forward human motion across frames similar to the mechanism of introducing noise in the forward process. We assume that the human motion sequence will eventually reach the initial distribution within total $N$ steps. (additional $N-f+1$ steps are necessary from $x_{f}$ to reach the initial distribution $x_{end}$). Consequently, we utilize a transformer-based diffusion model to sequentially produce the decoded features during the reverse process. The final human mesh sequence is returned by a mesh head using SMPL [33] human body model. The structure of DN (diffusion network) is illustrated in Fig. 6.
@@
The concatenation and separation strategy has substantially reduced the number of steps from $f\times N$ to $N$. However, it is important to note that this strategy does not effectively consider motion patterns across frames due to simple concatenation, which may result in temporal inconsistencies and non-smooth 3D mesh predictions.
@@ 3.3 DiffMesh
To address the previously mentioned limitations in our two baselines, we present DiffMesh, a novel solution for video-based HMR. The overall architecture is depicted in Fig. 5. Our framework takes input sequence with $f$ frames, with the objective of outputting a human mesh sequence consisting of $f$ frames. Additionally, we extract the conditional feature $c_{i}$ using the backbone, incorporating a conditional features generation block (more details can be found in the supplementary Sec. 2). In contrast to Baseline 1 and Baseline 2, DiffMesh employs a distinct approach that accounts for the forward and backward motion within the diffusion model.
Forward Process: We make the assumption that human motion consistently influences the human mesh vertices (in a latent space) across frames, denoted as $x_{i}$ to $x_{i+1}$, where $i\in\left\{1,\dots f-1\right\}$. This motion pattern can be conceptualized as a specific noise introduced to the mesh vertices feature of the preceding frame:
@@
where $m_{i}$ represents the motion pattern from $i_{th}$ frame to $(i+1)_{th}$ frame. Consequently, $x_{1}$ can be gradually diffused into $x_{f}$ after $f-1$ forward diffusion steps. Importantly, achieving $x_{f}$ does not denote the conclusion of the forward diffusion process. We assume that human motion will eventually reach the initial distribution $x_{end}$ through total $N$ diffusion steps (similar to the total $N$ steps to get the Gaussian distribution in the conventional diffusion model). More details and ablation study are provided in the supplementary Sec. 3. To arrive at this initial distribution $x_{end}$, DiffMesh requires $N-f+1$ additional forward diffusion steps from $x_{f}$. Therefore, the total number of steps remains $N$.
@@
Reverse Process: Our objective during the reverse process is to obtain output features $x^{\prime}_{1}\cdots x^{\prime}_{f}$. Initiating from the initial distribution $x_{end}$, we design a transformer-based diffusion model, as depicted in Fig. 5, to sequentially produce decoded features. After $N-f+1$ steps, our diffusion model yields $x^{\prime}_{f}$, which serves as the target feature for the final frame $f$. Iteratively, we generate the output features $x^{\prime}_{i}$ where $i\in\left\{f-1,\dots 1\right\}$ over $f-1$ steps. The total number of steps remains $N$. In contrast to Baseline 2, DiffMesh demonstrates the capability to decode specific motion patterns within the same number of steps, while generating more precise and smooth output (verified in Sec. 4.4).
@@
Figure 6: (a) Two-stream transformer network: Given the input feature $x^{\prime}_{i}$ and corresponding conditional feature $c_{i}$, the diffusion model generates the predicted motion $m_{i-1}$ and the predicted previous conditional feature $\hat{c}_{i-1}$. (b) The self-attention block in our diffusion model. (c) The cross-attention block in our diffusion model. The notation $\oplus$ represents element-wise addition.
Network Design: In contrast to conventional diffusion-based methods [15, 41, 40] for image synthesis tasks that often rely on UNet [42] or vanilla transformer [50] backbones, our approach introduces a two-stream transformer design deeply integrated within our framework, as illustrated in Fig. 6. While conventional diffusion-based methods [15, 41, 40] only estimate noise during steps, our two-stream network predicts the motion features and the previous conditional feature during each step as depicted in Fig. 6 (c). This architecture enhances information integration through a two-stream process. Initially, it captures mesh and condition dependencies separately using two self-attention blocks. Then, these captured dependencies are merged by a cross-attention block efficiently to decode motion features. An auxiliary loss is computed by measuring the difference (MSE loss) between the predicted $\hat{c}_{i-1}$ and the ground truth $c_{i-1}$, which contributes to training a better transformer-based network during the reverse process.
Given the input feature $x^{\prime}_{i}$ and corresponding conditional feature $c_{i}$ at each step, our network produces the predicted motion $m_{i-1}$ as well as the predicted previous conditional feature $\hat{c}_{i-1}$. The feature of initial distribution (mesh template) is denoted as $x_{end}$. Similar to Equation 5, the mesh features $x^{\prime}_{i-1}$ in DiffMesh can be computed as
$\small x^{\prime}_{i-1}=\frac{1}{\sqrt{\beta_{i}}}(x^{\prime}_{i}-\frac{1-\beta_{i}}{\sqrt{1-\hat{\beta_{i}}}}{\color[rgb]{1,.5,0}m_{i-1}})+\sigma_{t}{\color[rgb]{1,.5,0}x_{end}},$ (7)
@@
3.4 Overall Training and Loss Functions
Based on the assumption that human motion consistently impacts the human mesh across consecutive frames, we postulate the existence of a latent space where the motion pattern $m_{i}$ can be conceptualized as a specific noise introduced to the mesh vertices’ features of the preceding frame $x_{i}$, as depicted in Eq. 6. Our objective is to derive such motion patterns for each time interval within this specific latent space during the reverse steps. Ultimately, the final mesh sequence is generated based on the estimated mesh features $x_{i}$ through MLP layers and the SMPL model.
Thus, besides supervising the diffusion noise in the diffusion models, we also supervise the SMPL parameters of each frame ($\Theta_{i}$ and $\beta_{i}$) where $i\in\left\{1,\dots f\right\}$ with the ground-truth SMPL parameters using the $L2$ loss during training. Following [22, 5], the 3D joint coordinates are regressed by forwarding the SMPL parameters to the SMPL human body model. Furthermore, as mentioned in Sec. 3.3, Mean Squared Error (MSE) loss is applied between the conditional feature $c_{i}$ and the predicted conditional features $\hat{c_{i}}$. More detailed information is included in the supplementary Sec. 2.
4 Experiments